OpenStudy (anonymous):

lim pi/4 of (5 cos x) ^ tan x

OpenStudy (johnweldon1993):

\[\large \lim_{x\rightarrow \pi/4} (5cos(x))^{tan(x)}\] If we just plug in \(\large \pi/4\) into our equation we will get \[\large 5cos(\pi/4)^{tan(\pi/4)}\] We know \(\large cos(\pi/4) = \frac{1}{\sqrt{2}}\) and we know \(\large tan(\pi/4) = 1\) So what we have is \[\large 5(\frac{1}{\sqrt{2}})\] \[\large \frac{5}{\sqrt{2}}\]